The Andrews are selling their home. They want to keep $150,000 after paying a 7% commission to a realtor. For how much must they sell the house(to the nearest dollar)?
Found 2 solutions by rfer, advedantatutor:
Answer by rfer(16322)

150000*1.07=$160,500

Answer by advedantatutor(4)

After paying 7% commission they will get to keep 93% of the selling price.
93% = 150,000
100% = x
cross multiply
93 x x = 150,000 x 100
x = (150,000 x 100)/93
x = 161, 290.32
Selling price should be 161,290.32
Check :
Commission of 7% = 7% of 161,290.32
= 7/100 x 161,290.32 = 11,290.32.
161,290.32 - 11,290.32 = 150,000
The Andrews get 150,000.